3. Nuclear chemistry



1. Calculate the binding energy per nucleon in MeV for the iron isotope with an atomic number of 26 and a mass number of 56, 56Fe26 which has an atomic mass of 55.9349 amu.



2. Calculate:



a. The rate constant for the radioactive disintegration of cobalt-60, an isotope used in cancer therapy. 60Co27 decays with a half-life of 5.2 years to produce 60Ni28.



b. The fraction and the percentage of a sample of the 60Co27 isotope that will remain



after 15 years.



c. How long it takes for a sample of 60Co27 to disintegrate to the extent that only 2 % of the original amount remains?



d. The half-life of 216Po84 is 0.16 s. How long does it take to reduce a sample of it to the negligible amount of 0.000010 % of the original amount (1.0 x 10-7 times the original amount)?

1
Expert's answer
2021-11-04T00:32:08-0400

Binding Energy: The amount of energy released when a nucleus is formed from its components.• Mass defect ΔM = Z×Mp + N×Mn-MAZ= Atomic no., N= No. of neutrons M= Mass of proton, M= Mass of neutron,MA= Atomic Mass


Binding Energy = ΔMc= ΔM × 931.5 MeV


Binding Energy Per Nucleon = ΔMc= (ΔM ×931.5 MeV)/ (no. of protons + no. of neutrons)


Here ΔM can be calculated from above equation as:


Z=26, N=30, Mp=1.0078 u, MN= 1.0086 u.

ΔM = 26(1.0078) +30(1.0086)-55.9349ΔM = 0.5259 u


Binding Energy= ΔMc= 0.5259 × 931.5 MeV = 489.87 MeV


Binding Energy Per Nucleon = (489.87)/56 = 8.74 MeV